Approximately 14 million Americans are addicted to drugs and alcohol. The federal government estimates that these addicts cost the U.S. economy $300 billion in medical expenses and lost productivity. Despite the enormous potential market, many biotech companies have shied away from funding research and development (R&D) initiatives to find a cure for drug and alcohol addiction. Your firm – Drug Abuse Sciences (DAS) – is a notable exception. It has spent $160 million to date working on a cure, but is now at a crossroads. It can either abandon its program or invest another $35 million today. Unfortunately, the firm’s opportunity cost of funds is 7 percent and it will take another five years before final approval from the Federal Drug Administration is achieved and the product is actually sold. Expected (year-end) profits from selling the drug are presented in the accompanying table.

Year 1 Year 2 Year 3 Year 4 Year 5 Year 6 Year 7 Year 8 Year 9

$0 $0 $0 $0 $13,200,000 $16,300,000 $18,000,000 $20,500,000 $21,800,000

What is the net present value of the project?

Instruction: Enter your response rounded to the nearest penny (two decimal places). Use a negative sign (-) where appropriate. $

Should DAS continue with its plan to bring the drug to market, or should it abandon the project?
Answer #1

Solution :- Net present value (NPV) = Present value of cash inflows - Present value of cash outflow.

Present value of cash inflows = 13200000 / (1.07)5 + 16300000 / (1.07)6 + 18000000 / (1.07)7 + 20500000 / (1.07)8 + 21800000 / (1.07)9

= 13200000 / 1.4026 + 16300000 / 1.5007 + 18000000 / 1.6058 + 20500000 / 1.7182 + 21800000 / 1.8385

= 9411093.68 + 10861597.92 + 11209366.048 + 11931090.68 + 11857492.521

= $ 55270640.849 (approx)

Present value of cash outflow = $ 35000000. (Given in question, $ 35 Million i.e., $ 35000000).

Accordingly, Net present value (NPV) of project = 55270640.849 - 35000000

= $ 20270640.849 (Rounded off to $ 20270640.85)

Conclusion :- Net present value (NPV) of project = $ 20270640.85 (Rounded off). As the NPV of project is positive, therefore, DAS can run its project to make available drug in the market instead of abandoning the project.